OpenStudy (anonymous):

A person spins the pointer and is awarded the amount indicated by the pointer. Cicle wth half of the circle is $2 The other half has two amounts one fourth of it is $10 dollars. The other one fourth is $5 dollars. It costs $6 to play the game. Determine: The expectation of a person who plays the game. The fair price to play the game. Show work

OpenStudy (amistre64):

sum of (amount * probability) = expected value

OpenStudy (amistre64):

2(1/2) + 10(1/4) + 5(1/4) - 6(0)

OpenStudy (amistre64):

or maybe: -4(1/2) + 4(1/4) -1(1/4) if we assume winning values after paying 6 bucks

OpenStudy (amistre64):

the definition of fair game seems to be when the expected value is equal to the amount paid to play

OpenStudy (ganpat):

2(1/2) + 10(1/4) + 5(1/4) - 6(0) @amistre64 : 6 (0) ?? m confused ..

OpenStudy (amistre64):

it was a stray thought :) since the sum probs = 1 already, youve got nothing to lose with; but then i decided to adjust winnings based on amount cost to play

OpenStudy (ganpat):

u r investing 6$.. so 6 (x) ... x number of games .. what say ??

OpenStudy (amistre64):

1 + 2.5 + 1.25 = 4.75 as expected winnings to 6 dollars paid -2 + 1 -.25 = -1.25 the expected value is to loose 1.25; so the game aint fair

OpenStudy (amistre64):

6 - 4.75 = 1.25 in other words; each way i see it; you are out 1.25 no matter which method you use to determine expected value

OpenStudy (amistre64):

since the expected winning is 4.75; a fair game would cost 4.75

OpenStudy (amistre64):

site is acting a little off today ....